# Implemetation of a shooting technique for quantum optimal control on spin qudits

### 23 Mar 2026 v1

@@ C.3 Optimization routine: natural gradient descent
Consider Problem ($\mathscr{P}_{g}$). Instead of optimizing $|Pg|$ subject to the constraint $E(g)=U_{f}$, we simplify the problem by optimizing a functional $J(g)=\ell(E(g))$ penalizing the discrepancy to the target $U_{f}$. In other words, we look for a $g$ that is admissible for Problem ($\mathscr{P}_{g}$) but need not be time-optimal. However, this is done using a so-called natural gradient descent (see e.g. [27]), a variant of the Gauss-Newton method for optimization. Its update rule $g\mapsto\hat{g}$ reads
$\hat{g}=g-\rho\left[E^{\prime}\left(g\right)^{\dagger}E^{\prime}\left(g\right)+\alpha I\right]^{-1}\nabla J\left(g\right),\quad\text{where}\quad\nabla J\left(g\right)=E^{\prime}(g)^{\dagger}\nabla\ell(E(g)).$ (40)
In Section C.3.1 below we provide different ways to gain intuition on this update rule. Via the Tikhonov regularization parameter $\alpha$, this method allows for a natural way to penalize the variation of the euclidean norm of $g$ and hence of the time $T=|Pg|$. Put differently, this means that we do not optimize the time directly but rather attempt to find an admissible point $g$ for which $|Pg|$ is just as large as it needs to be for reaching the target.
We use an optimize-then-discretize approach, also known as indirect method. This means that we first compute a descent direction such as $\nabla J(g)$ and only then introduce a numerical scheme to discretize the equations of motion. This is contrast to the use of an automatic differentiation engine which is part of the discretize-then-optimize framework, also known as direct method.
@@ C.3.1 Principle of natural gradient descent
Up to choosing an orthonormal basis $(g_{i})_{1\leq i\leq n}$ of $\operatorname{i}\mathfrak{su}(d)$, we may identify $g$ with its coordinates $\xi_{i}=\langle g,g_{i}\rangle$. Accordingly we may view $E$ as a map from $\mathbb{R}^{n}\to\mathbb{C}^{d\times d}$.
@@
The intuition behind splitting $J$ as $\ell$ composed with the model or end-point mapping $E$ is the following: when $\ell$ is convex, optimizing in the $U$ space is in general far easier than optimizing $J$ directly in the $g$ space due to non-convexity. To exploit this idea, the natural gradient step takes into account the model parametrization $g\mapsto U(1)=E(g)$. It projects $\nabla\ell(U(1))$, the ”good” descent direction in $U$ space, onto the range of the linearization of $E$ at $g$. Roughly speaking, the range of $E^{\prime}(g)$ describes the directions of motion locally available around $U(1)=E(g)$ within the state set $S=\{E(g),\,g\in\operatorname{i}\mathfrak{su}(d)\}$. This projection can be seen as preconditioning $\nabla J(g)$, the ”bad” gradient in $g$ space, by the regularized Gram matrix $E^{\prime}(g)^{\dagger}E^{\prime}(g)$ which describes how the parameters or degrees of freedom $g$ offer an ”independent” description of the state set $S$ around a the state $E(g)$.
@@
(i) Algorithm to calculate JVP: $E^{\prime}(g)\delta g$.
@@
Given $g$, we solve forward from $t=0$ to $t=1$ the augmented system $(U,\delta U)$ starting at $(I,0)$ where $\delta U(t)=-\operatorname{i}U(t)w(t)$ is the linearization of the trajectory $U$ in the direction $\delta g$ and $w$ is defined by Equation (31). The value at time $t=1$ of $\delta U$ is the desired quantity $E^{\prime}(g)\delta g$.
(ii) Algorithm to calculate VJP: $E^{\prime}(g)^{\dagger}S$.
@@
Given $g$, we solve for the terminal state $U(1)$ and evolve backward from $t=1$ to $t=0$ the augmented system $x=(U,v,z)$ starting at $x_{1}(1)=U(1)$, $x_{2}(1)=\operatorname{i}x_{1}^{\dagger}S$ and $x_{3}(1)=0$. The equation of motion for $x_{1}$ and $x_{2}=v$ are respectively given by (30) and (33) with minus signs to account for backward motion, and that of $x_{3}=z$ is $\dot{z}=-\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v)h_{j}$. In this way the value at time $t=0$ of $z$ is the desired quantity $R=E^{\prime}(g)^{\dagger}S$.
(iii) Algorithm to calculate $\nabla J(g)$:
Given $g$, apply (ii) with $S=\nabla\ell(E(g))$.
(iv) Algorithm to calculate combined VJP and JVP: $E^{\prime}(g)^{\dagger}E^{\prime}(z)\delta g$.
@@
Find the update step of the natural gradient descent requires solving a symmetric positive-definite linear system involving the gram matrix of $E^{\prime}(g)$ regularized by $\alpha I$. Since $\alpha>0$ and to avoid creating the huge matrix $E^{\prime}(g)$ of order scaling like $d^{2}$, we appeal to matrix-free methods of Krylov-type. For instance GMRES [32, 35].
@@
Given $\delta g$, we apply (i) to get $S:=E^{\prime}(g)\delta g$ and then (ii) to obtain $R=E^{\prime}(g)^{\dagger}S$. Then $R+\alpha\delta g$ is the desired image of $\delta g$ under the linear map defining the normal equations. The step-size $\rho$ is calculated using the golden-section method, a gradient-free optimization method for unimodal functions [30]. This assumption is usually not verified but the method empirically works decently well in our setting.
(v) Initialization.
@@
Whenever $g^{(0)}$ is orthogonal to $\mathscr{H}_{0}$, the linear span of the control Hamiltonians, it is an equilibrium point of the optimization procedure. In particular for some targets $U_{f}$, the reasonable initial guess $g^{(0)}=\operatorname{i}\log(U_{f})$ could be one such equilibrium point.
@@
Accordingly a random initial-guess is preferable without further a priori knowledge. For the purpose of minimizing $T=|Pg|$ subject to $E(g)=U_{f}$, it is additionally preferable for $|Pg^{(0)}|$ to be realtively ”small” and to have the penalty $\alpha$ be quite large.
@@
Typically $\alpha=10^{-3}$ is a robust choice with respect to the choice of target $U_{f}$ and dimension $d$. If $\alpha$ is larger, we expect $|Pg|$ to be smaller when the algorithm stops, but this may take a lot more additional iterations to occur.
(vi) Validation.
@@
Our stopping criterion is of the based on the value of the objective function and of the form $J(g)\leq\tau$ where $\epsilon$ is a given absolute tolerance. Once the algorithm stops, we validate the candidate numerical solution $g$ using an independent and high-fidelity ODE solver to compute the $U(1)$ associated to this $g$. We evaluate the corresponding discrepancy to the target $J_{\rm val}(g)$ and check whether it is indeed less then or equal to $\tau$. Typically $\tau=10^{-4}$ if $\ell$ is the infidelity $\ell(x)=1-|\operatorname{Tr}(UU_{f}^{\dagger})|/d$.
@@
Given a uniform grid of the interval $[0,1]$ with stepsize $\delta t$, we optimize until the stopping criterion $J(g)\leq\tau$ is met. If the result is not validated in the sense of (vii), that is $J_{\rm val}(g)>\tau$, then we refine the mesh. For instance by a factor two $\delta t_{\rm new}=\delta t/2$. Keeping $g$ as initial guess, we then repeat the previous steps until $J_{\rm val}\leq\tau$ or the number of iterations exceeds a pre-defined threshold. This can be seen as a continuation method over the mesh-size.

### 10 Apr 2026 v2

@@ C.2.1 The adjoint state method
C.2.2 Proofs of the JVP and VJP expressions
Proof of Lemma C.4 (JVP).
We denote by $\delta U(t)$ the first order variation of the trajectory $U(t)$ solution to (46) in response to a perturbation $\delta M$ of $M$
$\delta U(t)=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}U_{M+\varepsilon\delta M}(t)\right|_{\varepsilon=0}$ (56)
By definition $E^{\prime}(M)\delta M=\delta U(1)$. In Equation (46), keep in mind that $U$ itself depends on $M$. Then differentiating with respect to $M$ the initial condition and both sides of the differential equation yields that the first variation of $U$ is the solution starting at $\delta U(0)=0$ to $\dot{\delta U}=-\operatorname{i}U(A+B)$ where
$A$ $=\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left((\delta U)^{\dagger}H_{j}UM+U^{\dagger}H_{j}\delta UM+h_{j}\delta M\right)h_{j},$ (57)
$B$ $=U^{\dagger}H\delta U.$ (58)
We can simplify the $A$ term by observing that $\operatorname{Re}\operatorname{Tr}((\delta U)^{\dagger}H_{j}UM)=\operatorname{Re}\operatorname{Tr}(U^{\dagger}H_{j}\delta UM)$, which means that the two first contributions from the left are equal. Then multiplying $\delta U$ by the inverse propagator $U^{\dagger}$ will cancel the contribution of the $B$ term. Indeed if $w(t)=\operatorname{i}U(t)^{\dagger}\delta U(t)$ then $\dot{w}=\operatorname{i}\dot{U}^{\dagger}\delta U+(A+B)$ and, given the differential equation on $U$, the first term simplifies to $\operatorname{i}\dot{U}^{\dagger}\delta U=-B$. Altogether we see that $\dot{w}=A$. Substituting $\delta U=-\operatorname{i}Uw$ in the $A$ terms yields the Equation (47) on $w$ because $w(0)=0$ since $\delta U(0)=0$. ∎
Proof of Lemma C.5 (VJP).
The aim is to find the Hermitian traceless matrix $R$ such that
$\operatorname{Re}\operatorname{Tr}\left(S^{\dagger}\delta U(1)\right)=\operatorname{Re}\operatorname{Tr}\left(R\delta M\right)$ (59)
for any perturbation $\delta M$ of $M$. Define $w=\operatorname{i}U\delta U$ as in Lemma C.4. Observe that for any function $v$
$\operatorname{Re}\operatorname{Tr}\left(v(1)^{\dagger}w(1)\right)=\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left(\dot{v}^{\dagger}w+v^{\dagger}\dot{w}\right)\mathrm{d}t$ (60)
by the fundamental theorem of calculus. Using the differential equation on $w$ lets us rewrite Equation (60) as
$\operatorname{Re}\operatorname{Tr}\left(v(1)^{\dagger}w(1)\right)$ $-\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left((\dot{v}+2\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(v^{\dagger}h_{j})h_{j}M)^{\dagger}w\right)\mathrm{d}t$ (61)
$=$ $\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left(\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(v^{\dagger}h_{j})h_{j}\delta M\right)\mathrm{d}t$ (62)
Notice the right-hand side only involves the perturbations $\delta M$ but not $w$. This means that choosing $v$ as the solution to Equation (49), the left-hand side becomes equal to the left-hand side of (59). Then choosing $R$ as in Eq. (48) gives (49) and concludes the proof. ∎
@@ C.3 Optimization routine: natural gradient descent
Consider Problem ($\mathscr{P}_{M}$). Instead of optimizing $|P(M)|$ subject to the constraint $E(M)=U_{f}$, we simplify the problem by optimizing a functional $J(M)=\ell(E(M))$ penalizing the discrepancy to the target $U_{f}$. In other words, we look for a $M$ that is admissible for Problem ($\mathscr{P}_{M}$) but need not be time-optimal. However, this is done using a so-called natural gradient descent (see e.g. [30]), a variant of the Gauss-Newton method for optimization. Its update rule $M\mapsto\hat{M}$ reads
$\hat{M}=M-\rho\left[E^{\prime}\left(M\right)^{\dagger}E^{\prime}\left(M\right)+\alpha I\right]^{-1}\nabla J\left(M\right),\quad\text{where}\quad\nabla J\left(M\right)=E^{\prime}(M)^{\dagger}\nabla\ell(E(M)).$ (63)
In Section C.3.1 below we provide different ways to gain intuition on this update rule. Via the Tikhonov regularization parameter $\alpha$, this method allows for a natural way to penalize the variation of the Frobenius (Hilbert-Schmidt) norm of $M$ and hence of the time $T=|P(M)|$. Put differently, this means that we do not optimize the time directly but rather attempt to find an admissible point $M$ for which $|P(M)|$ is just as large as it needs to be for reaching the target.
We use an optimize-then-discretize approach, also known as indirect method. This means that we first compute a descent direction such as $\nabla J(M)$ and only then introduce a numerical scheme to discretize the equations of motion. This is contrast to the use of an automatic differentiation engine which is part of the discretize-then-optimize framework, also known as direct method.
@@ C.3.1 Principle of natural gradient descent
Up to choosing an orthonormal basis $(M_{i})_{1\leq i\leq n}$ of $\operatorname{i}\mathfrak{su}(d)$, we may identify $M$ with its coordinates $\xi_{i}=\langle M,M_{i}\rangle$. Accordingly we may view $E$ as a map from $\mathbb{R}^{n}\to\mathbb{C}^{d\times d}$.
@@
The intuition behind splitting $J$ as $\ell$ composed with the model or end-point mapping $E$ is the following: when $\ell$ is convex, optimizing in the $U$ space is in general far easier than optimizing $J$ directly in the $M$ space due to non-convexity. To exploit this idea, the natural gradient step takes into account the model parametrization $M\mapsto U(1)=E(M)$. It projects $\nabla\ell(U(1))$, the ”good” descent direction in $U$ space, onto the range of the linearization of $E$ at $M$. Roughly speaking, the range of $E^{\prime}(M)$ describes the directions of motion locally available around $U(1)=E(M)$ within the state set $S=\{E(M),\,M\in\operatorname{i}\mathfrak{su}(d)\}$. This projection can be seen as preconditioning $\nabla J(M)$, the ”bad” gradient in $M$ space, by the regularized Gram matrix $E^{\prime}(M)^{\dagger}E^{\prime}(M)$ which describes how the parameters or degrees of freedom $M$ offer an ”independent” description of the state set $S$ around the state $E(M)$.
@@
Algorithm to calculate JVP: $E^{\prime}(M)\delta M$.
@@
Given $M$, we solve forward from $t=0$ to $t=1$ the augmented system $(U,\delta U)$ starting at $(I,0)$ where $\delta U(t)=-\operatorname{i}U(t)w(t)$ is the linearization of the trajectory $U$ in the direction $\delta M$ and $w$ is defined by Equation (47). The value at time $t=1$ of $\delta U$ is the desired quantity $E^{\prime}(M)\delta M$.
Algorithm to calculate VJP: $E^{\prime}(M)^{\dagger}S$.
@@
Given $M$, we solve for the terminal state $U(1)$ and evolve backward from $t=1$ to $t=0$ the augmented system $x=(U,v,z)$ starting at $x_{1}(1)=U(1)$, $x_{2}(1)=\operatorname{i}x_{1}^{\dagger}S$ and $x_{3}(1)=0$. The equation of motion for $x_{1}$ and $x_{2}=v$ are respectively given by (46) and (49) with minus signs to account for backward motion, and that of $x_{3}=z$ is $\dot{z}=-\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v)h_{j}$. In this way the value at time $t=0$ of $z$ is the desired quantity $R=E^{\prime}(M)^{\dagger}S$.
Algorithm to calculate $\nabla J(M)$:
Given $M$, apply (ii) with $S=\nabla\ell(E(M))$.
Algorithm to calculate combined VJP and JVP: $E^{\prime}(M)^{\dagger}E^{\prime}(z)\delta M$.
@@
Find the update step of the natural gradient descent requires solving a symmetric positive-definite linear system involving the gram matrix of $E^{\prime}(M)$ regularized by $\alpha I$. Since $\alpha>0$ and to avoid creating the large non-sparse matrix $E^{\prime}(M)$ of order scaling like $d^{2}$, we appeal to matrix-free methods of Krylov-type. For instance GMRES [34, 38].
@@
Given $\delta M$, we apply (a) to get $S:=E^{\prime}(M)\delta M$ and then (b) to obtain $R=E^{\prime}(M)^{\dagger}S$. Then $R+\alpha\delta M$ is the desired image of $\delta M$ under the linear map defining the normal equations. The step-size $\rho$ is calculated using the golden-section method, a gradient-free optimization method for unimodal functions [33]. This assumption is usually not verified but the method empirically works decently well in our setting.
Initialization.
@@
Whenever $M^{(0)}$ is orthogonal to the linear span of the control Hamiltonians, it is an equilibrium point of the optimization procedure. In particular for some targets $U_{f}$, the reasonable initial guess $M^{(0)}=\operatorname{i}\log(U_{f})$ could be one such equilibrium point.
@@
Accordingly a random initial-guess is preferable without further a priori knowledge. For the purpose of minimizing $T=|P(M)|$ subject to $E(M)=U_{f}$, it is additionally preferable for $|P(M^{(0)})|$ to be relatively ”small” and to choose the penalty $\alpha$ quite large.
@@
Typically $\alpha=10^{-3}$ is a robust choice with respect to the choice of target $U_{f}$ and dimension $d$. If $\alpha$ is larger, we expect $|P(M)|$ to be smaller when the algorithm stops, but this may require a lot more additional iterations.
Validation.
@@
Our stopping criterion is based on the value of the objective function and of the form $J(M)\leq\tau$ where $\epsilon$ is a given absolute tolerance. Once the algorithm stops, we validate the candidate numerical solution $M$ using an independent and high-fidelity ODE solver to compute the $U(1)$ associated to this $M$. We evaluate the corresponding discrepancy to the target $J_{\rm val}(M)$ and check whether it is indeed less then or equal to $\tau$. Typically $\tau=10^{-4}$ if $\ell$ is the infidelity $\ell(x)=1-|\operatorname{Tr}(UU_{f}^{\dagger})|/d$.
@@
Given a uniform grid of the interval $[0,1]$ with stepsize $\delta t$, we optimize until the stopping criterion $J(M)\leq\tau$ is met. If the result is not validated in the sense of (f), that is $J_{\rm val}(M)>\tau$, then we refine the mesh. For instance by a factor two $\delta t_{\rm new}=\delta t/2$. Keeping $M$ as initial guess, we then repeat the previous steps until $J_{\rm val}\leq\tau$ or the number of iterations exceeds a pre-defined threshold. This can be seen as a continuation method over the mesh-size.